Integrated Nanosystems Research Facility

The HF room in the BION will be closed today as we will be doing maintenance on the Dicing Saw.

Due to the flood in the front office the contractors will be dry-walling and painting all day today.

Tomorrow morning at approximately 8am facilities will be adding a new N2, lab air and vacuum line to the room for the dicing saw. We don’t expect there to be any downtime but just pay attention when working in there to make sure everything is working properly. Ill post again if there are any updates.

The E-Beam 1 is now back up. The Trion and Mark 50 are still down.

INRF Staff

The INRF cleanroom will be closed March 18th – 28th in order to replace the buildings air compressor. As you know, we have been experiencing problems with low pressure and oil in the air lines for years now. This should remedy those issues and open up more possibilities for the cleanroom in the future.

The BION will remain open during this time.

NOTE: All compressed air to the Engineering Gateway West wing building will be shut down during this time. This will effect all 3rd floor labs as well.


Location: Cal(IT)2 (BION)
Disruption/Outage: DI System
Dates: Thursday, December 17th, 2015 & Friday, December 18th, 2015
Duration: Scheduled to begin at 7:00am on Thursday, 12-17-2015 and ending by 6:00pm on Friday, 12-18-2015
Details: This outage is to perform sanitization of the DI systems throughout Cal(IT)2.

Location: Engineering Gateway (INRF)
Disruption/Outage: DI System
Dates: Monday, December 14th, 2015 & Tuesday, December 15th, 2015
Duration: Scheduled to begin at 7:00am on Monday, 12-14-2015 and ending by 6:00pm on Tuesday, 12-15-2015
Details: This outage is to perform sanitization of the DI systems throughout Engineering Gateway.

We appreciate your cooperation and apologize for any inconvenience this may cause.  If you have any questions or concerns, please contact Tammy Kelleher, at 949-824-8715 or After Hours contact the Central Plant at 949-824-5520


STS is down for repair on some of the ceramics. (We are working with STS to get the parts ASAP)

Ebeam1 is down for maintenance to regen the cryo. Should be back up tomorrow.

The pump for the Trion needs to be sent out for maintenance. The Trion will be down until further notice.

Facilities will be performing maintenance on exhaust fans 7 & 9 this Saturday from 6am-2pm. The back-up fans should handle load while the maintenance is performed but please pay close attention to warning lights and anything unusual if you are working in the BION during this time.

This Friday, Praxair will be connecting a temporary nitrogen tank trailer in order to keep the lines active during Mondays removal/install. This will impact visitor parking in the Calit2 loading lock (the loading dock will still be accessible). Jake will start shutting down most equipment starting Thursday night and then bring them back up once the install is complete. The cleanrooms will remain open during this time.